How do you solve 5x-4(2x+11)=13 ?
Answer by gonzo(654)

since you don't have a parentheses around (5x-4), i'm assuming that the equation is:
5x - (4*(2x + 11)) = 13 (original equation)
removing parentheses, this becomes:
5x - 8x - 44 = 13
combine like terms;
-3x - 44 = 13
add 44 to both sides:
-3x = 57
divide both sides by -3:
x = -57/3 = -19
-----
substitute -19 for x in original equation:
5*(-19) - 4(2*(-19) + 11) = 13
remove parentheses:
5*(-19) - 8*(-19) - 44 = 13
combine like terms:
-3(-19) - 44 = 13
perform indicated operations:
57 - 44 = 13
13 = 13
-----
answer is:
x = -19
-----